# The Dark Dimension, the Swampland, and the Dark Matter Fraction Composed of Primordial Near-Extremal Black Holes

### v1

@@ Abstract
In a recent publication we studied the decay rate of primordial black holes perceiving the dark dimension, an innovative five-dimensional (5D) scenario that has a compact space with characteristic length-scale in the micron range. We demonstrated that the rate of Hawking radiation of 5D black holes slows down compared to 4D black holes of the same mass. Armed with our findings we showed that for a species scale of ${\cal O}(10^{9}~{}{\rm GeV})$, an all-dark-matter interpretation in terms of primordial black holes should be feasible for black hole masses in the range $10^{14}\lesssim M/{\rm g}\lesssim 10^{21}$. As a natural outgrowth of our recent study, herein we calculate the Hawking evaporation of near-extremal 5D black holes. Using generic entropy arguments we demonstrate that Hawking evaporation of higher-dimensional near-extremal black holes proceeds at a slower rate than the corresponding Schwarzschild black holes of the same mass. Assisted by this result we show that if there were 5D primordial near-extremal black holes in nature, then a PBH all-dark-matter interpretation would be possible in the mass range $10^{7}\sqrt{\beta}\lesssim M/{\rm g}\lesssim 10^{21}$, where $\beta$ is a parameter that controls the difference between mass and charge of the associated near-extremal black hole.
@@ I Introduction
As demonstrated in Montero:2022prj , applying the AdS-DC to dS space could help elucidate the radiative stability of the cosmological hierarchy $\Lambda/M_{p}^{4}\sim 10^{-120}$, because it connects the size of the compact space $R_{\perp}$ to the dark energy scale $\Lambda^{-1/4}$ via $R_{\perp}\sim\lambda\ \Lambda^{-1/4}$, where the proportionality factor is estimated to be within the range $10^{-1}<\lambda<10^{-4}$. Actually, the previous relation between $R_{\perp}$ and $\Lambda$ derives from constraints by theory and experiment. On the one hand, since the associated Kaluza-Klein (KK) tower contains massive spin-2 bosons, the Higuchi bound Higuchi:1986py provides an absolute upper limit to $\alpha$, whereas explicit string calculations of the vacuum energy (see e.g. Itoyama:1986ei ; Itoyama:1987rc ; Antoniadis:1991kh ; Bonnefoy:2018tcp ) yield a lower bound on $\alpha$. All in all, the theoretical constraints lead to $21/4\leq\alpha\leq 1/$ 2. On the other hand, experimental arguments (e.g. constraints on deviations from Newton’s gravitational inverse-square law Lee:2020zjt and neutron star heating Hannestad:2003yd ) lead to the conclusion encapsulated in $R_{\perp}\sim\lambda\ \Lambda^{-1/4}$; namely, that there is one extra dimension of radius $R_{\perp}$ in the micron range, and that the lower bound for $\alpha=1/4$ is basically saturated Montero:2022prj . A theoretical amendment on the connection between the cosmological and KK mass scales confirms $\alpha=1/4$ Anchordoqui:2023laz . Assembling all this together, we can conclude that the KK tower of the new (dark) dimension opens up at the mass scale $m_{\rm KK}\sim 1/R_{\perp}$. For the dark dimensions scenario, the five-dimensional (5D) Planck scale (or species scale where gravity becomes strong Dvali:2007hz ; Dvali:2007wp ; Cribiori:2022nke ; vandeHeisteeg:2023dlw ) is given by
@@
where $M_{p}$ is the reduced Planck mass. Thus, since the size of dark dimension in the micron scale, $m_{\rm KK}\sim 1~{}{\rm eV}$ and so $M_{*}\sim 10^{9}~{}{\rm GeV}$.
In previous work Anchordoqui:2022txe ; Anchordoqui:2022tgp , we first calculated the decay rate of primordial black holes (PBH) perceiving the dark dimension and demonstrated that the rate of Hawking radiation slows down compared to 4D black holes of the same mass. Then, we used this result to show that the mass range supporting a 5D PBH all-dark-matter interpretation is extended compared to that in the 4D theory by 3 orders of magnitude in the low mass region. As a natural outgrowth of this work, herein we study the Hawking evaporation of near-extremal 5D black holes. More concretely, we generalize the 4D results obtained in deFreitasPacheco:2023hpb to $d$-dimensions. We then discuss the impact of our findings in assessing the dark matter fraction that could be composed of PBHs. Since Hawking evaporation of near-extremal 5D black holes proceeds at a slower rate than the corresponding Schwarzschild black holes of the same mass, we show herein that near extremality could further relax the lower mass bound range of a PBH all-dark-matter interpretation.
@@ II Primordial black hole dark matter interpretation
It has long been speculated that black holes could be produced from the collapse of large amplitude fluctuations in the early universe Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj . For an order of magnitude estimate of the black hole mass $M$, we first note that the cosmological energy density scales with time $t$ as $\rho\sim 1/(Gt^{2})$ and the density needed for a region of mass $M$ to collapse within its Schwarzschild radius is $\rho\sim c^{6}/(G^{3}M^{2})$, so that primordial black holes (PBHs) would initially have around the cosmological horizon mass Carr:2020xqk
@@
Nevertheless, an all-dark-matter interpretation in terms of PBHs is severely constrained by observations Carr:2020xqk ; Green:2020jor ; Villanueva-Domingo:2021spv . Of relevance to our investigation, the extragalactic $\gamma$-ray background Carr:2009jm and on the CMB spectrum Clark:2016nst constrain PBH evaporation of black holes with masses $\lesssim 10^{17}~{}{\rm g}$, whereas the non-observation of microlensing events from the MACHO Macho:2000nvd , EROS EROS-2:2006ryy , Kepler Griest:2013aaa , Icarus Oguri:2017ock , OGLE Niikura:2019kqi and Subaru-HSC Croon:2020ouk collaborations constrain black holes with masses $\gtrsim 10^{21}~{}{\rm g}$.
@@
$T_{s}\sim\frac{1}{2\pi r_{s}}\sim\sqrt{\frac{3}{64}}\ \frac{1}{\pi}\ \frac{M_{p}\ \Lambda^{1/8}}{\lambda^{1/2}\ M^{1/2}}\sim\left(\frac{M}{10^{10}~{}{\rm g}}\right)^{-1/2}~{}{\rm MeV}\,,$ (4)
@@
$r_{s}(M)\sim\frac{1}{M_{*}}\left[\frac{16}{3}\ \frac{M}{M_{*}}\ \zeta(4)\right]^{1/2}$ (5)
is the 5D Schwarzschild radius Myers:1986un and $\zeta(4)=\pi^{2}/90$. The second relation in (4) applies to the dark dimension scenario. For the numerical estimate of (4), we have taken $\lambda\sim 10^{-3}$ as suggested by astrophysical observations Anchordoqui:2022ejw ; Noble:2023mfw . It is evident that 5D black holes are colder than 4D black holes of the same mass. The Hawking radiation causes a 4D black hole to lose mass at the following rate Keith:2021guq
@@
Integrating (7) we can parametrized the 5D black hole lifetime as a function of its mass and temperature,
$\tau_{s}\sim 14\ \Bigg{(}\frac{M}{10^{11.33}~{}{\rm g}}\Bigg{)}^{2}\ \left(\frac{4}{\sum_{i}c_{i}(T_{s})\ \tilde{f}\ \Gamma_{s}}\right)^{-1}~{}{\rm Gyr}\,,$ (8)
where we have used (4) to estimate that $T_{s}\sim$ and therefore $c_{i}(T_{s})$ receives a contribution of 6 from neutrinos and 2 from photons, yielding $\sum_{i}c_{i}(T_{s})\ \tilde{f}\ \Gamma_{s}=4$. Armed with (8) it is easily seen that for a species scale of ${\cal O}(10^{9}~{}{\rm GeV})$, an all-dark-matter interpretation in terms of 5D black holes must be feasible for masses in the range
@@
As we advanced in the Introduction, this range is extended compared to that in the 4D theory by 3 orders of magnitude in the low mass region.
@@ III Near-extremal black holes
where $q$ is the integer-quantized electric charge of the particle and $M_{p}$ is the reduced Planck mass Arkani-Hamed:2006emk . Setting $e=e^{\prime}=\sqrt{4\pi\alpha}$ the (4D) Schwinger effect together with the WGC lead to a bound on the minimum black hole mass of near extremal black holes with evaporation time longer than the age of Universe, $M_{ne}\gtrsim 5\times 10^{15}~{}{\rm g}(m_{e^{\prime}}/10^{9}~{}{\rm GeV})^{-2}$ Bai:2019zcd .
@@ IV Higher-dimensional Schwinger pair production
[…]
@@
where $E^{\prime}$ is the dark electric field, $e^{\prime}Q=\sqrt{4\pi}M/M_{*}^{(d-2)/2}$, and $J$ is the spin of the produced particles Bachas:1992bh . A point worth noting at this juncture is that for $d>6$ the RN-dS solution is gravitational unstable Konoplya:2008au ; Konoplya:2013sba and so we focus our calculation on the interesting case of $d=5$ that characterizes the dark dimension scenario Montero:2022prj . For $d=5$, the spin $J$ is half-integer and so (14) can be rewritten as
@@
$r_{+,5d}=\left[\frac{M+\sqrt{M^{2}-(e^{\prime}Q)^{2}\ M_{*}^{3}/(4\pi)}}{M_{*}^{3}}\right]^{1/2}=\sqrt{\frac{M}{M_{*}^{3}}}\,,$ (16)
and that of a 4D black hole $r_{+,4d}=M/M_{p}^{2}$ with the same $M$ and $e^{\prime}Q$. It follows that $r_{+,5d}>r_{+,4d}\Leftrightarrow M<M_{p}^{4}/M_{*}^{3}$, which if we take $M_{*}\sim 10^{9}~{}{\rm GeV}$ implies that $M<10^{45}~{}{\rm GeV}$ and $r_{+,5d}<1~{}\mu{\rm m}$. This in turn entails that for the length scale of interest, the outer horizon of a 5D RN black hole is larger than the corresponding 4D black hole. If this were the case, then the electric field strength in the outer horizon would be smaller and it would easier to suppress Schwinger production pairs in five than in four dimensions.
@@ V Higher-dimensional near-extremal black hole decay rate
Next, in line with our stated plan, we investigate how near-extremal black holes could modify the PBH range given in Eq. (9). To do so, we consider a black hole with $M\sim 10^{5}~{}{\rm g}$. From (4) we see that such a black hole has a temperature $T_{s}\sim 707~{}{\rm MeV}$. This means that $c_{i}(T_{s})$ receives a contribution of 2 from photons, 6 from neutrinos, 4 from electrons, 4 from muons, 12 from quarks (up, down, and strange), and 8 from gluons, yielding $\sum_{i}c_{i}(T_{s})\ \tilde{f}\ \Gamma_{s}=53/3$. Substituting these figures into (8) we find that the lifetime of a $10^{5}~{}{\rm g}$ Schwarzschild black hole is $\tau_{s}\sim 4\times 10^{-5}~{}{\rm yr}$. For a near extremal black hole of the same mass, the temperature would be $T_{ne}\sim 10^{-6}\sqrt{\beta}~{}{\rm eV}$, where we have used (23). Bearing this in mind we find that the lifetime of the near-extremal black hole would be $\tau_{ne}\sim 20/\sqrt{\beta}~{}{\rm Gyr}$. Now, a rough order of magnitude estimate suggests that if there were 5D primordial near-extremal black holes in nature, then a PBH all-dark-matter interpretation would be possible in the mass range
$10^{7}\sqrt{\beta}\lesssim M/{\rm g}\lesssim 10^{21}\,.$ (25)

### v2

@@ Abstract
In a recent publication we studied the decay rate of primordial black holes perceiving the dark dimension, an innovative five-dimensional (5D) scenario that has a compact space with characteristic length-scale in the micron range. We demonstrated that the rate of Hawking radiation of 5D black holes slows down compared to 4D black holes of the same mass. Armed with our findings we showed that for a species scale of ${\cal O}(10^{10}~{}{\rm GeV})$, an all-dark-matter interpretation in terms of primordial black holes should be feasible for black hole masses in the range $10^{14}\lesssim M/{\rm g}\lesssim 10^{21}$. As a natural outgrowth of our recent study, herein we calculate the Hawking evaporation of near-extremal 5D black holes. Using generic entropy arguments we demonstrate that Hawking evaporation of higher-dimensional near-extremal black holes proceeds at a slower rate than the corresponding Schwarzschild black holes of the same mass. Assisted by this result we show that if there were 5D primordial near-extremal black holes in nature, then a primordial black hole all-dark-matter interpretation would be possible in the mass range $10^{5}\sqrt{\beta}\lesssim M/{\rm g}\lesssim 10^{21}$, where $\beta$ is a parameter that controls the difference between mass and charge of the associated near-extremal black hole.
@@ I Introduction
As demonstrated in Montero:2022prj , applying the AdS-DC to dS space could help elucidate the origin of the cosmological hierarchy $\Lambda/M_{p}^{4}\sim 10^{-120}$, because it connects the size of the compact space $R_{\perp}$ to the dark energy scale $\Lambda^{-1/4}$ via $R_{\perp}\sim\lambda\ \Lambda^{-1/4}$, where the proportionality factor is estimated to be within the range $10^{-1}<\lambda<10^{-4}$. Actually, the previous relation between $R_{\perp}$ and $\Lambda$ derives from constraints by theory and experiment. On the one hand, since the associated Kaluza-Klein (KK) tower contains massive spin-2 bosons, the Higuchi bound Higuchi:1986py provides an absolute upper limit to $\alpha$, whereas explicit string calculations of the vacuum energy (see e.g. Itoyama:1986ei ; Itoyama:1987rc ; Antoniadis:1991kh ; Bonnefoy:2018tcp ) yield a lower bound on $\alpha$. All in all, the theoretical constraints lead to $21/4\leq\alpha\leq 1/$ 2. On the other hand, experimental arguments (e.g. constraints on deviations from Newton’s gravitational inverse-square law Lee:2020zjt and neutron star heating Hannestad:2003yd ) lead to the conclusion encapsulated in $R_{\perp}\sim\lambda\ \Lambda^{-1/4}$; namely, that there is one extra dimension of radius $R_{\perp}$ is in the micron range, and that the lower bound for $\alpha=1/4$ is basically saturated Montero:2022prj . A theoretical amendment on the connection between the cosmological and KK mass scales confirms $\alpha=1/4$ Anchordoqui:2023laz . Assembling all this together, we can conclude that the KK tower of the new (dark) dimension opens up at the mass scale $m_{\rm KK}\sim 1/R_{\perp}$. For the dark dimensions scenario, the five-dimensional (5D) Planck scale (or species scale where gravity becomes strong Dvali:2007hz ; Dvali:2007wp ; Cribiori:2022nke ; vandeHeisteeg:2023dlw ) is given by
@@
where $M_{p}$ is the reduced Planck mass. Thus, since the size of dark dimension in the micron scale, $m_{\rm KK}\sim 1~{}{\rm eV}$ and so $10^{9}\lesssim M_{*}/{\rm GeV}\lesssim 10^{10}$.
Early universe phenomena and the nature of dark matter are among the most strategic science cases of theoretical high energy physics. As potentially the first density perturbations to collapse during the early universe, primordial black holes (PBHs) provide our earliest landmarks to probe the very early universe, at energies between the QCD phase transition and the Planck scale. The corresponding energy scales are out of the reach of existing cosmological probes. Much of the parameter space characterizing the PBH abundance has been constrained by existing probes, but a large window remains open, where PBHs around asteroid mass ($10^{-15}$ to $10^{-10}M_{\odot}$) could make up the entirety of dark matter Carr:2020xqk ; Green:2020jor ; Villanueva-Domingo:2021spv ; LISACosmologyWorkingGroup:2023njw . The detection of PBHs could provide a cornerstone for our perception of the physics processes in the very early universe. This significant reward motivates new investigations on this subject.
In previous work Anchordoqui:2022txe ; Anchordoqui:2022tgp , we first calculated the decay rate of PBHs perceiving the dark dimension and demonstrated that the rate of Hawking radiation slows down compared to 4D black holes of the same mass. Then, we used this result to show that the mass range supporting a 5D PBH all-dark-matter interpretation is extended compared to that in the 4D theory by 3 orders of magnitude in the low mass region. As a natural outgrowth of this work, herein we study the Hawking evaporation of near-extremal 5D black holes. More concretely, we generalize the 4D results obtained in deFreitasPacheco:2023hpb to $d$-dimensions. We then discuss the impact of our findings in assessing the dark matter fraction that could be composed of PBHs. Since Hawking evaporation of near-extremal 5D black holes proceeds at a slower rate than the corresponding Schwarzschild black holes of the same mass, we show herein that near extremality could further relax the lower mass bound range of a PBH all-dark-matter interpretation.
@@ II Primordial black hole dark matter interpretation
It has long been speculated that black holes could be produced from the collapse of large amplitude fluctuations in the early universe Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx ; Carr:1975qj . For an order of magnitude estimate of the black hole mass $M$, we first note that the cosmological energy density scales with time $t$ as $\rho\sim 1/(Gt^{2})$ and the density needed for a region of mass $M$ to collapse within its Schwarzschild radius is $\rho\sim c^{6}/(G^{3}M^{2})$, so that PBHs would initially have around the cosmological horizon mass Carr:2020xqk
@@
Nevertheless, an all-dark-matter interpretation in terms of PBHs is severely constrained by observations Carr:2020xqk ; Green:2020jor ; Villanueva-Domingo:2021spv ; LISACosmologyWorkingGroup:2023njw . Of relevance to our investigation, the extragalactic $\gamma$-ray background Carr:2009jm and the spectrum of the cosmic microwave background (CMB) Clark:2016nst constrain PBH evaporation of black holes with masses $\lesssim 10^{17}~{}{\rm g}$, whereas the non-observation of microlensing events from the MACHO Macho:2000nvd , EROS EROS-2:2006ryy , Kepler Griest:2013aaa , Icarus Oguri:2017ock , OGLE Niikura:2019kqi and Subaru-HSC Croon:2020ouk collaborations constrain black holes with masses $\gtrsim 10^{21}~{}{\rm g}$.
@@
$T_{s}\sim\frac{1}{r_{s}}\sim\left(\frac{M}{10^{12}~{}{\rm g}}\right)^{-1/2}~{}{\rm MeV}\,,$ (4)
@@
$r_{s}(M)\sim\frac{1}{M_{*}}\left[\frac{2}{3\,\pi}\ \frac{M}{M_{*}}\right]^{1/2}$ (5)
is the 5D Schwarzschild radius Myers:1986un . The numerical estimate of (4) applies to the dark dimension scenario with $M_{*}\sim 10^{10}~{}{\rm GeV}$, which is consistent with astrophysical observations Anchordoqui:2022ejw ; Noble:2023mfw . [^1] We have taken the highest possible value of $M_{*}$ to remain conservative in the estimated bound on $f_{\rm PBH}$. It is evident that 5D black holes are colder than 4D black holes of the same mass. The Hawking radiation causes a 4D black hole to lose mass at the following rate Keith:2021guq
@@
Integrating (7) we can parametrize the 5D black hole lifetime as a function of its mass and temperature,
$\tau_{s}\sim 13.8\ \Bigg{(}\frac{M}{10^{12}~{}{\rm g}}\Bigg{)}^{2}\ \left(\frac{6}{\sum_{i}c_{i}(T_{s})\ \tilde{f}\ \Gamma_{s}}\right)~{}{\rm Gyr}\,,$ (8)
where we have used (4) to estimate that $T_{s}\sim 1~{}{\rm MeV}$ and therefore $c_{i}(T_{s})$ receives a contribution of 6 from neutrinos, 4 for electrons, and 2 from photons, yielding $\sum_{i}c_{i}(T_{s})\ \tilde{f}\ \Gamma_{s}=6$. Armed with (8) we can estimate the bound on the 5D PBH abundance by a simple rescaling procedure of the $d=4$ bounds on the fraction of dark matter composed of primordial black holes $f_{\rm PBH}$. The key point for such a rescaling is that for a given photon energy, or equivalently a given Hawking temperature, we expect a comparable limit on $f_{\rm PBH}$ for both $d=4$ and $d=5$. For example, from (3) and (4) we see that the constraint of $f_{\rm PBH}\lesssim 10^{-3}$ for 4D black holes with $M_{\rm BH}\sim 10^{16}~{}{\rm g}$, should be roughly the same for the abundance of 5D black holes with $M_{\rm BH}\sim 10^{12}~{}{\rm g}$. Now, since in $d=4$ for $M_{\rm BH}\sim 10^{17}~{}{\rm g}$ we have $f_{\rm PBH}\sim 1$, this implies the same abundance for 5D black holes of $M_{\rm BH}\sim 10^{14}~{}{\rm g}$. By duplicating this procedure for heavier black holes we conclude that for a species scale of ${\cal O}(10^{10}~{}{\rm GeV})$, an all-dark-matter interpretation in terms of 5D black holes must be feasible for masses in the range
@@
This range is extended compared to that in the 4D theory by 3 orders of magnitude in the low mass region.
@@ III Near-extremal black holes
where $q$ is the integer-quantized electric charge of the particle and $M_{p}$ is the reduced Planck mass Arkani-Hamed:2006emk ; Harlow:2022ich . Setting $e=e^{\prime}=\sqrt{4\pi\alpha}$ the (4D) Schwinger effect together with the WGC lead to a bound on the minimum black hole mass of near extremal black holes with evaporation time longer than the age of Universe, $M_{ne}\gtrsim 5\times 10^{15}~{}{\rm g}(m_{e^{\prime}}/10^{9}~{}{\rm GeV})^{-2}$ Bai:2019zcd .
@@ IV Higher-dimensional Schwinger pair production
Before proceeding, we pause to note that the FL inequality (11) remains the same in any number of dimensions, since the gauge coupling has units of Energy2-d/2 Montero:2021otb . However, it is important to stress that the FL bound only applies to black holes of size comparable to the cosmological horizon and therefore it is not of direct interest for scales smaller than $R_{\perp}$. The dS-WGC is relevant to black holes with a horizon radius smaller than $R_{\perp}$ Antoniadis:2020xso . However, for large $\ell_{4}$ values, dS-WGC constraints on the particle spectrum can also be safely neglected. Hereafter, we proceed under the assumption of a (nearly) flat 5D Minkowski background and neglect the last term in (13). We further assume that (10) is satisfied. For details, see the Appendix.
@@
where $E^{\prime}$ is the dark electric field, $e^{\prime}Q\to\sqrt{4\pi}M/M_{*}^{(d-2)/2}$, and $J$ is the spin of the produced particles Bachas:1992bh . Throughout, the arrow indicates we are considering near-extremal rather than extremal black holes. A point worth noting at this juncture is that for $d>6$ the RN-dS solution is gravitationally unstable Konoplya:2008au ; Konoplya:2013sba and so we focus our calculation on the interesting case of $d=5$ that characterizes the dark dimension scenario Montero:2022prj . For $d=5$, the spin $J$ is half-integer and so (14) can be rewritten as
@@
$r_{+,5d}=\left[\frac{M+\sqrt{M^{2}-(e^{\prime}Q)^{2}\ M_{*}^{3}/(4\pi)}}{M_{*}^{3}}\right]^{1/2}\to\sqrt{\frac{M}{M_{*}^{3}}}\,,$ (16)
and that of a 4D black hole $r_{+,4d}\to M/M_{p}^{2}$ with the same $M$ and $e^{\prime}Q$. It follows that $r_{+,5d}>r_{+,4d}\Leftrightarrow M<M_{p}^{4}/M_{*}^{3}$, which if we take $M_{*}\sim 10^{9}~{}{\rm GeV}$ implies that $M<10^{45}~{}{\rm GeV}$ and $r_{+,5d}<1~{}\mu{\rm m}$. This in turn entails that for the length scale of interest, the outer horizon of a 5D RN black hole is larger than the corresponding 4D black hole. If this were the case, then the electric field strength in the outer horizon would be smaller and it would be easier to suppress Schwinger production pairs in five than in four dimensions.
@@ V Higher-dimensional near-extremal black hole decay rate
Next, in line with our stated plan, we investigate how near-extremal black holes could modify the PBH range given in Eq. (9). To do so, we consider a black hole with $M\sim 10^{5}~{}{\rm g}$. From (4) we see that such a black hole has a temperature $T_{s}\sim 4~{}{\rm GeV}$. This means that $c_{i}(T_{s})$ receives a contribution of 2 from photons, 6 from neutrinos, 12 from charged leptons (electrons, muons, and taus), 48 from quarks (up, down, strange, and charm), and 24 from gluons, yielding $\sum_{i}c_{i}(T_{s})\ \tilde{f}\ \Gamma_{s}=45$. Substituting these figures into (6) we find that the lifetime of a $10^{5}~{}{\rm g}$ Schwarzschild black hole is $\tau_{s}\sim 10^{-5}~{}{\rm yr}$. For a near extremal black hole of the same mass, the temperature would be $T_{ne}\sim 10^{-5}\sqrt{\beta}~{}{\rm eV}$, where we have used (23). Bearing this in mind we find that the lifetime of the near-extremal black hole would be $\tau_{ne}\sim 15/\sqrt{\beta}~{}{\rm Gyr}$. Now, the temperature of the near-extremal black hole is below the CMB temperature and hence there are no constraints from electromagnetic signals. The bound simply comes from the black hole survival probability. Then, a rough order of magnitude estimate suggests that if there were 5D primordial near-extremal black holes in nature, then a PBH all-dark-matter interpretation would be possible in the mass range
$10^{5}\sqrt{\beta}\lesssim M/{\rm g}\lesssim 10^{21}\,.$ (25)